# Quasiparticle interference and spectral function of the UTe2 superconductive surface band

### v1

@@ B.4.1 Non-chiral order parameters
Under $m_{x}$ and $C_{2x}$, $G(E,k_{x},k_{\#},x)$ transforms respectively as
$G(E,k_{x},k_{\#},x)=\rho^{-1}_{\mathcal{C}}(m_{x})G(-k_{x},k_{\#},x)\rho_{\mathcal{C}}(m_{x})$ (49)
$G(E,k_{x},k_{\#},x)=\rho^{-1}_{\mathcal{C}}(C_{2x})G(k_{x},-k_{\#},-x)\rho_{\mathcal{C}}(C_{2x})$ (50)
@@
$G(E,k_{x},k_{\#},x)=\rho^{-1}_{\mathcal{C}}(\mathcal{T})G(E,-k_{x},-k_{\#},x)\rho_{\mathcal{C}}(\mathcal{T})$ (51)
@@ B.4.2 Chiral order parameters
We first discuss the symmetries of $G_{i}$ for chiral order parameters when $E\neq 0$. Chiral order parameters respect only a subset of the symmetries of non-chiral order parameters. We can separate the chiral order parameters that are symmetric or not under $m_{x}$. When they are symmetric under $m_{x}$, $\mathcal{T}m_{x}$ is broken. This is the case for $A_{u}+iB_{3u}$ and $B_{1u}+iB_{2u}$. In contrast, when $m_{x}$ is broken, $\mathcal{T}m_{x}$ remains a symmetry as for $A_{u}+iB_{1u}$, $B_{2u}+iB_{3u}$, $A_{u}+iB_{2u}$ and $B_{1u}+iB_{3u}$. Following the discussion of non-chiral order parameters, those respecting $m_{x}$ are mirror symmetric along the $k_{x}$ axis, while order parameters respecting $\mathcal{T}m_{x}$ show a mirror symmetry along the $k_{\#}$ axis.
@@
$S_{x}$circ $\times$ $\times$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$S_{y}$circ $\times$ $\circ$ $\times$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$S_{z}$circ $\circ$ $\times$ $\times$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
@@ B.5 Node structure for $B_{2u}$ and $B_{3u}$
Figure 19: Bulk spectral function in the (0-11)-plane for non-chiral pairing $B_{2u}$ and $B_{3u}$ at $E=0$. The projection of the nodes onto the surface are indicated by the black circles. We take $\Delta_{0}=0.3$ meV and $\eta=10$ $\mu$eV.
@@ B.6.1 Spin-polarization of $G_{b}$
[…]
$\left[G_{\mathrm{BdG}}(E,\mathbf{k})\right]_{ee}=\left[E-h(\mathbf{k})-\Delta(\mathbf{k})\left(E+h^{*}(-\mathbf{k})\right)\Delta^{\dagger}(\mathbf{k})\right]^{-1}$ (54)
Because $h(\mathbf{k})$ is diagonal in the spin basis and $\Delta(\mathbf{k})$ and $\Delta^{\dagger}(\mathbf{k})$ have opposite spin polarizations, Eq. (54) shows that $G_{b}(E,\mathbf{k}_{\parallel})$ is not spin-polarized. Consequently, any spin polarization in $G_{s}$ originates exclusively from $G_{i}$. In the following we note $G_{\Delta}=\left[G_{\mathrm{BdG}}\right]_{eh}$ and $G_{\Delta^{\dagger}}=\left[G_{\mathrm{BdG}}\right]_{he}$. $G_{\Delta}$ have the polarization of $\Delta$ and $G_{\Delta^{\dagger}}$ of $\Delta^{\dagger}$.
@@ B.6.2 Spin-polarization of $G_{i}$
$T(E,\mathbf{k}_{\parallel})=\begin{pmatrix}T_{e}(E,\mathbf{k}_{\parallel})&T_{\Delta}(E,\mathbf{k}_{\parallel})\\ T_{\Delta^{\dagger}}(E,\mathbf{k}_{\parallel})&T_{h}(E,\mathbf{k}_{\parallel})\\ \end{pmatrix}$ (55)
@@
$G(E,\mathbf{k},x)=\begin{pmatrix}G_{e}^{x}(E,\mathbf{k}_{\parallel})&G_{\Delta}^{x}(E,\mathbf{k}_{\parallel})\\ G_{\Delta^{\dagger}}(E,\mathbf{k}_{\parallel})&G_{h}^{x}(E,\mathbf{k}_{\parallel})\\ \end{pmatrix}$ (56)
@@
$\left[G_{i}\right]_{ee}=G^{x=1}_{e}T_{e}G^{x=-1}_{e}+G^{x=1}_{e}T_{\Delta}G^{x=-1}_{\Delta^{\dagger}}$ (57)
$+G^{x=1}_{\Delta}T_{\Delta^{\dagger}}G^{x=-1}_{e}+G^{x=1}_{\Delta}T_{h}G^{x=-1}_{\Delta^{\dagger}}$ (58)
@@
$\vec{d}(\mathbf{k}).\vec{\sigma}=\sum\limits_{i}d_{i}(\mathbf{k})\sigma_{i}$ $G^{x}_{\Delta}\sim\sum\limits_{i}\alpha^{x}_{i}(k_{x},k_{\#})\sigma_{i}$ (59)
@@
$G_{\Delta}^{x=1}T_{h}G^{x=-1}_{\Delta^{\dagger}}\sim\sum\limits_{ij}\alpha_{i}^{x=1}\alpha_{i}^{x=-1}\sigma_{i}\sigma_{j}$ (60)
@@
As an example, let’s study the $B_{3u}$ order parameter with $d$-vector
$\vec{d}_{B_{3u}}(\mathbf{k})=\Delta_{0}\left(0,\sin(k_{z}),\sin(k_{y})\right)^{T}$ (61)
Since $\vec{d}_{B_{3u}}$ lacks a $k_{x}$-dependence, the spin polarization along the $x$-direction will exhibit mirror symmetry with respect to the $x$-axis. Due to time-reversal symmetry $\mathcal{T}$, which flips the spin, the polarization will be anti-mirror symmetric along the $k_{\#}$-axis. This anti-mirror symmetry arises because the combined operation $\mathcal{T}m_{x}$ enforces the corresponding spatial transformation. The same reasoning can be extended to the other order parameters and explains the numerical findings.
@@ Appendix C Additional plots at $E=0$
In Sec. IV we have presented the results for the surface spectral function, the spin-resolved surface spectral function along $\widehat{x}$-axis and the QPI pattern at zero-energy for various SC pairings. For the sake of completeness, in this Appendix we provide also the bulk spectral function, the spin-resolved surface spectral functions along $\widehat{y}$-axis and $\widehat{z}$-axis, and the JDOS pattern at $E=0$.
@@ C.1 Spin-resolved spectral function along $\widehat{y}$- and $\widehat{z}$-axis
The spin-resolved surface spectral function along $\widehat{y}$-axis and $\widehat{z}$-axis are given in Figs. 20 and 22 for non-chiral pairings, and in Figs. 21 and 23 for chiral pairings. We emphasize that the spin-resolved surface spectral function cancels for $B_{1u}$ and $B_{3u}$ pairings along $\widehat{y}$-axis, and for $B_{2u}$ and $B_{3u}$ pairings along $\widehat{z}$-axis.
[…]
@@ C.2 Joint density of states
The JDOS provides an intuitive picture of possible peak locations in the QPI pattern. In Figs. 24 and  25 we shown the JDOS at $E=0$. The observed peaks correspond to wavevectors connecting two regions of high intensity in the surface spectral function. From the bulk and edge contribution to the surface spectral function, shown respectively in Figs. 6, 7 and Figs. 8, 9, we observe that the spectral function tends to show similar regions of high intensity for all the OPs. This implies that the peaks in the autocorrelation spectrum will share roughly the same wavevectors. This shows that the autocorrelation spectrum does not provide sufficient information to differentiate between different OPs for this system. However the experimental measurements should not be compared with the JDOS but with the results of the full T-matrix calculations (the QPI patterns). In the full calculations due to phase cancellations the intensity and the position of the peaks can be very different from one OP to the next, thus the comparison with the experiments allow us to identify the correct OP.
@@ Appendix D Study of energy dependence
D.1 Surface spectral function at $E\neq 0$
Figures 26 and 27 depict the evolution of the surface spectral function for an energy varying from 0 to 0.25 meV. We note that, as also discussed in the main text, the amplitude of the surface spectral function is highly dependent on energy, thus superconducting surface states visible at low energy can disappear at higher energy for some pairings, or inversely, superconducting surface states absent at low energy can appear at higher energy for others.
We also note that at $E\neq 0$ the mirror-symmetry with respect to the $k_{x}$-axis and $k_{\#}$-axis is preserved for non-chiral pairings, but can be broken for some of the chiral pairings. The explanation for this is that, although the triplet order parameters associated to $A_{u}$, $B_{1u}$, $B_{2u}$ and $B_{3u}$ pairings are not invariant under the spatial symmetries of the group $Immm$, they remain invariant under symmetries in the superconducting phase, because a gauge transformation can remove a $U(1)$ phase [54]. However, when two order parameters that transform differently under a crystalline symmetry $g$ are combined, such as in $B_{1u}+iB_{2u}$ for example, gauge invariance alone is insufficient to restore the crystalline symmetry $g$. The symmetry breaking for the chiral order parameters is reflected in the surface spectral function plots: as shown in Figs. 26 and 27, the edge contribution to the surface spectral function for $E\neq 0$ remains mirror-symmetric for the non-chiral pairings, while this symmetry is broken in the chiral case.
D.2 QPI patterns at $E\neq 0$
Figues 28 and 29 describe the evolution of the QPI pattern for an energy varying from 0 to 0.25 meV. Note that the ${\bf q}_{1}$ peak visible for $B_{3u}$, $A_{u}+iB_{1u}$, $A_{u}+iB_{2u}$, $A_{u}+iB_{3u}$, $B_{1u}+iB_{3u}$ and $B_{2u}+iB_{3u}$ pairings is stable in position and remains visible up to $E\approx 0.15$ meV.

### v2

@@ B.4.1 Non-chiral order parameters
Under $m_{x}$ and $C_{2x}$, $G(E,k_{x},k_{\#},x)$ transforms, respectively, as
$G(E,k_{x},k_{\#},x)=\rho^{-1}_{\mathcal{C}}(m_{x})G(-k_{x},k_{\#},x)\rho_{\mathcal{C}}(m_{x})~,$ (51)
$G(E,k_{x},k_{\#},x)=\rho^{-1}_{\mathcal{C}}(C_{2x})G(k_{x},-k_{\#},-x)\rho_{\mathcal{C}}(C_{2x})~.$ (52)
@@
$G(E,k_{x},k_{\#},x)=\rho^{-1}_{\mathcal{C}}(\mathcal{T})G(E,-k_{x},-k_{\#},x)\rho_{\mathcal{C}}(\mathcal{T})~.$ (53)
@@ B.4.2 Chiral order parameters
We first discuss the symmetries of $G_{i}$ for chiral order parameters when $E\neq 0$. Chiral order parameters respect only a subset of the symmetries of non-chiral order parameters. We can separate the chiral order parameters that are symmetric or not under $m_{x}$. When they are symmetric under $m_{x}$, $\mathcal{T}m_{x}$ is broken. This is the case for $A_{u}+iB_{3u}$ and $B_{1u}+iB_{2u}$. In contrast, when $m_{x}$ is broken, $\mathcal{T}m_{x}$ remains a symmetry for $A_{u}+iB_{1u}$, $B_{2u}+iB_{3u}$, $A_{u}+iB_{2u}$ and $B_{1u}+iB_{3u}$. Following the discussion of non-chiral order parameters, those respecting $m_{x}$ are mirror symmetric along the $k_{x}$ axis, while order parameters respecting $\mathcal{T}m_{x}$ show a mirror symmetry along the $k_{\#}$ axis.
@@
$S_{x}$ $\circ$ $\times$ $\times$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$S_{y}$ $\circ$ $\times$ $\circ$ $\times$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$S_{z}$ $\circ$ $\circ$ $\times$ $\times$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
@@ B.5 Node structure for $B_{2u}$ and $B_{3u}$
[…]
@@ B.6.1 Spin-polarization of $G_{b}$
$\left[G_{b}(E,\mathbf{k}_{\parallel})\right]_{ee}=\int dk_{\perp}\left[G_{\mathrm{BdG}}(E,k_{\perp},\mathbf{k}_{\parallel})\right]_{ee}~,$ (55)
$\left[G_{\mathrm{BdG}}(E,\mathbf{k})\right]_{ee}=\bigg[E-h(\mathbf{k})$
$-\Delta(\mathbf{k})\left[E+h^{*}(-\mathbf{k})\right]^{-1}\Delta^{\dagger}(\mathbf{k})\bigg]^{-1}~.$ (56)
Because $h(\mathbf{k})$ is diagonal in the spin basis and $\Delta(\mathbf{k})$ and $\Delta^{\dagger}(\mathbf{k})$ have opposite spin polarizations, Eq. (B.6.1) shows that $G_{b}(E,\mathbf{k}_{\parallel})$ is not spin-polarized. Consequently, any spin polarization in $G_{s}$ originates exclusively from $G_{i}$. In the following, we note $G_{\Delta}=\left[G_{\mathrm{BdG}}\right]_{eh}$ and $G_{\Delta^{\dagger}}=\left[G_{\mathrm{BdG}}\right]_{he}$. $G_{\Delta}$ have the polarization of $\Delta$ and $G_{\Delta^{\dagger}}$ of $\Delta^{\dagger}$.
@@ B.6.2 Spin-polarization of $G_{i}$
$T(E,\mathbf{k}_{\parallel})=\begin{pmatrix}T_{e}(E,\mathbf{k}_{\parallel})&T_{\Delta}(E,\mathbf{k}_{\parallel})\\ T_{\Delta^{\dagger}}(E,\mathbf{k}_{\parallel})&T_{h}(E,\mathbf{k}_{\parallel})\\ \end{pmatrix}~$. (57)
@@
$G(E,\mathbf{k},x)=\begin{pmatrix}G_{e}^{x}(E,\mathbf{k}_{\parallel})&G_{\Delta}^{x}(E,\mathbf{k}_{\parallel})\\ G_{\Delta^{\dagger}}(E,\mathbf{k}_{\parallel})&G_{h}^{x}(E,\mathbf{k}_{\parallel})\\ \end{pmatrix}~$. (58)
@@
$\left[G_{i}\right]_{ee}=G^{x=1}_{e}T_{e}G^{x=-1}_{e}+G^{x=1}_{e}T_{\Delta}G^{x=-1}_{\Delta^{\dagger}}$
$+G^{x=1}_{\Delta}T_{\Delta^{\dagger}}G^{x=-1}_{e}+G^{x=1}_{\Delta}T_{h}G^{x=-1}_{\Delta^{\dagger}}~.$ (59)
@@
$\vec{d}(\mathbf{k})\cdot\vec{\sigma}=\sum\limits_{i}d_{i}(\mathbf{k})\sigma_{i}~,$ $G^{x}_{\Delta}\sim\sum\limits_{i}\alpha^{x}_{i}(k_{x},k_{\#})\sigma_{i}~,$ (60)
@@
$G_{\Delta}^{x=1}T_{h}G^{x=-1}_{\Delta^{\dagger}}\sim\sum\limits_{ij}\alpha_{i}^{x=1}\alpha_{i}^{x=-1}\sigma_{i}\sigma_{j}~.$ (61)
@@
As an example, let us study the $B_{3u}$ order parameter with $d$-vector expression given in Eq. (13). In the simplest case considered in this work ${\bf d}_{B_{3u}}$ lacks a $k_{x}$-dependence, and the spin polarization along the $x$-direction will exhibit mirror symmetry with respect to the $x$-axis. Due to time-reversal symmetry $\mathcal{T}$, which flips the spin, the polarization will be anti-mirror symmetric along the $k_{\#}$-axis. This anti-mirror symmetry arises because the combined operation $\mathcal{T}m_{x}$ enforces the corresponding spatial transformation. More terms respecting all symmetries can be added to the $B_{3u}$ $d$-vector. The general method stays the same and the symmetry of the spin polarization can be inferred from the product of specific components of the $d$-vector. The same reasoning can also be extended to the other order parameters and explains the numerical findings.
@@ Appendix C Additional plots at $E=0$
In Sec. IV we presented the results for the surface spectral function, the spin-resolved surface spectral function along $\widehat{x}-$axis and the QPI pattern at zero-energy for various SC pairings. For the sake of completeness, in this Appendix we also provide the bulk spectral function, the spin-resolved surface spectral functions along $\widehat{y}-$axis and $\widehat{z}-$axis, and the JDOS pattern at $E=0$.
@@ C.1 Spin-resolved spectral function along $\widehat{y}-$ and $\widehat{z}-$axis
The spin-resolved surface spectral function along $\widehat{y}-$axis and $\widehat{z}-$axis are given in Figs. 22 and 24 for non-chiral pairings, and in Figs. 23 and 25 for chiral pairings. We emphasize that the spin-resolved surface spectral function cancels for $B_{1u}$ and $B_{3u}$ pairings along $\widehat{y}-$axis, and for $B_{2u}$ and $B_{3u}$ pairings along $\widehat{z}-$axis.
Non-chiral order parameters are time-reversal symmetric which imposes that surface states with opposite momenta have opposite spin polarizations, resulting in the absence of backscattering from non-magnetic impurities [hofmann2013theory]. Specifically, two states related by time-reversal symmetry satisfy $|u(\mathbf{k})\rangle=\mathcal{T}|v(-\mathbf{k})\rangle$, with $\mathcal{T}^{\,2}=-1$ in spinful systems. This implies that surface states with opposite momenta will not contribute to the QPI pattern. For chiral order parameters, time-reversal symmetry $\mathcal{T}$ is broken, so this argument does not apply directly. However, chiral order parameters still exhibit non-trivial magnetic symmetries combined with crystalline symmetries, such as $\mathcal{T}m_{x}$ for $A_{u}+iB_{1u}$ which impose constraints on the spin-resolved spectral function.
@@ C.2 Joint density of states
The JDOS provides an intuitive picture of possible peak locations in the QPI pattern. In Figs. 26 and  27 we shown the JDOS at $E=0$. The observed peaks correspond to wavevectors connecting two regions of high intensity in the surface spectral function. From the bulk and edge contribution to the surface spectral function, shown respectively in Figs. 6, 7 and Figs. 8, 9, we observe that the spectral function tends to show similar regions of high intensity for all the OPs. This implies that the peaks in the autocorrelation spectrum will share roughly the same wavevectors. This shows that the autocorrelation spectrum does not provide sufficient information to differentiate between different OPs for this system. However, the experimental measurements should not be compared with the JDOS but with the results of the full T-matrix calculations (the QPI patterns). In the full calculations due to phase cancellations the intensity and the position of the peaks can be very different from one OP to the next, thus the comparison with the experiments allow us to identify the correct OP.
@@ Appendix D Study of energy dependence
D.1 Surface energy band
Figures 29 and 30 show the surface energy band along a path ABCDA on the (0-11)-plane depicted in Fig. 28, at $\Delta_{0}=0.3$ meV and $\eta=0.03$ meV. This path is chosen to highlight the main characteristics, in particular the high intensity features in the surface energy band. We took a smaller value for the damping ($\eta=\Delta_{0}/10$) compared with the one used for the surface spectral function and QPI ($\eta=\Delta_{0}/3$) to better see the dispersing bands.
Figure 28: Path ABCDA in the (0-11)-plane.
Figure 29: Energy band dispersion (in meV) along the path ABCDA shown in Fig. 28 for non-chiral pairings at $\Delta_{0}=0.3$ meV and $\eta=0.03$ meV.
Figure 30: Same as in Fig. 29 for chiral pairings.
D.2 Surface spectral function at $E\neq 0$
Figures 31 and 32 depict the evolution of the surface spectral function for an energy varying from 0 to 0.25 meV. We note that, as also discussed in the main text, the amplitude of the surface spectral function is highly dependent on energy, thus superconducting surface states visible at low energy can disappear at higher energy for some pairings, or inversely, superconducting surface states absent at low energy can appear at higher energy for others.
We also note that at $E\neq 0$ the mirror-symmetry with respect to the $k_{x}$-axis and $k_{\#}$-axis is preserved for non-chiral pairings, but can be broken for some of the chiral pairings. The explanation for this is that, although the triplet order parameters associated with $A_{u}$, $B_{1u}$, $B_{2u}$ and $B_{3u}$ pairings are not invariant under the spatial symmetries of the group $Immm$, they remain invariant under symmetries in the superconducting phase, because a gauge transformation can remove a $U(1)$ phase [ono2020refined]. However, when two order parameters that transform differently under a crystalline symmetry $g$ are combined, such as in $B_{1u}+iB_{2u}$ for example, gauge invariance alone is insufficient to restore the crystalline symmetry $g$. The symmetry breaking for the chiral order parameters is reflected in the surface spectral function plots: as shown in Figs. 31 and 32, the edge contribution to the surface spectral function for $E\neq 0$ remains mirror-symmetric for the non-chiral pairings, while this symmetry is broken in the chiral case.
D.3 QPI patterns at $E\neq 0$
Figures 33 and 34 describe the evolution of the QPI pattern for an energy varying from 0 to 0.25 meV. Note that the ${\bf q}_{1}$ peak visible for $B_{3u}$, $A_{u}+iB_{1u}$, $A_{u}+iB_{2u}$, $A_{u}+iB_{3u}$, $B_{1u}+iB_{3u}$ and $B_{2u}+iB_{3u}$ pairings is stable in position and remains visible up to $E\approx 0.15$ meV.
Figure 31: Surface spectral function for non-chiral pairings at different energies. We take $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV.
Figure 32: Same as in Fig. 31 for chiral pairings.
Figure 33: QPI pattern for non-chiral pairings at different energies. For better visibility here we restrict the plots to the first Brillouin zone. We take $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV.
Figure 34: Same as in Fig. 33 for chiral pairings.